a rectangular shaped swimming pool is 12.5meters long and 10.5meters wide when my sister swims from one corner to the other corner diagonally opposite , how many meters to the nearest tenth of a meters does she swim ? . (please with solutions )
Answer by mananth(16946)

Pythagoras theorem states that
The diagonal and the sides of the pool form a right triangle
Hypotenuse ^2= leg1^2+leg2^2

leg1 = 12.5 m
Leg2^2= 10.5 m

Diagonal ^2= 12.5 ^2 + 10.5 ^2
Diagonal ^2= 156.25 + 110.25
Diagonal ^2= 266.5
take the square root
Diagonal= 16.3 m
she swims 16.3 m